# A compact model of Escherichia coli core and biosynthetic metabolism

### v2

@@ A.5.5 Conversion to apparent turnover numbers
The above procedure produces (after conversion back to a linear scale) a set of adjusted turnover parameters, $\mathbf{k}_{i}^{*}$ and scalings, $\boldsymbol{\theta}^{*}$. In order to retrieve the typical enzyme efficiencies and condition-specific scaling factors introduced in 25, we simply parametrise the solution by factorising the scaling terms $\theta_{j}$ as
$\sigma_{j}^{*}\equiv\bar{\sigma}^{*}~{}\tau_{j}^{*}$ (36)
Here, $\bar{\sigma}^{*}$ is the geometric mean of the scalings across conditions, which we interpret as typical enzyme saturation level across conditions, while $\tau_{j}^{*}$ is a residual scaling factor fluctuating around $1$ which required to account for differences in total measured enzymes across conditions. The typical enzyme efficiencies ($\kappa_{i}$) introduced in 25 are then simply recovered by incorporating the $\bar{\sigma}^{*}$ constant into the fitted turnover parameters:
@@ A.5.6 Potential extensions
We conclude this section by noting that the procedure described above assumes that the original parameter set – the set of $k_{\rm cat}$ values as proxies for apparent $k_{\rm cat}$ values in section A.5.3 – is sufficiently good to produce a realistic flux distribution to use as a reference for the adjustment step. If this is not the case, then a multi-start approach can be implemented, where multiple turnover parameter sets are first generated by perturbing the original parameter set, and each is used to generate a separate set of flux distributions. Each reference set is then provided as an input to problem (35), and the solution achieving the lowest objective is chosen in the end. The perturbed parameter set may be generated either randomly (for example, by introducing log-normal noise in the original turnover parameter vector) or systematically. The latter could be achieved, for example, by identifying reactions with zero-predicted flux but high measured abundance of the associated enzyme. By systematically increasing the corresponding turnover number, one can ”encourage” these reactions to be included in the reference set, potentially leading to the exploration of more relevant reference distributions than achieved by random perturbation.
@@ Appendix B Supplementary Figures
Supplementary Figure 1: Biosynthesis of amino acids from core metabolism precursors in iCH360. The Escher maps (including Suppl. Fig. 2 to 4) are deposited along with the code.
@@
Supplementary Figure 2: Biosynthesis of pyrimidine and purine (deoxy)ribonucleotides from in iCH360.
@@
Supplementary Figure 3: Biosynthesis of saturated and unsaturated fatty acids from acetyl-CoA. The map for the saturated fatty acids is taken from Escher [king_escher_2015].
Supplementary Figure 4: Metabolism of one-carbon compounds in iCH360.
@@
Supplementary Figure 5: Comparison of model size between ECC, ECC2, iCH360 and iML1515. To allow for a fair comparison, pseudo-reactions (e.g. exchange reactions) were excluded from the count. Note that gene annotation were not available in the SBML model of ECC2 accompanying its publication[hadicke_ecolicore2_2017].
@@
Supplementary Figure 6: Maximal biomass fluxes achieved by iCH360 and the parent model iML1515 for aerobic and anaerobic growth across multiple carbon sources. In all cases, the substrate uptake flux was bounded to $10$ mmol/gDW/h.
@@
Supplementary Figure 7: Comparison of production envelopes between iCH360 and the parent model iML1515. Top: production of ethanol, acetate and succinate underaerobic growth on lactate. Bottom: production of ethanol, acetate, lactate and succinate under aerobic growth on glycerol. Note, that the dashed line representing the production envelope of iML1515 is sometimes hidden behind the blue line corresponding to iCH360.
@@
Supplementary Figure 8: Fitness loss associated with network disruptions classified as complete disruptions (disruption of all catalytic edges for a reaction) and full primary disruption (disruption of all primary catalytic associations, but with remaining secondary ones) are not significantly different (Wilcoxon rank-sum test, $p=0.063$).
@@
Supplementary Figure 9: Proteome allocation predictions across conditions obtained with the turnover parameter set from [heckmann_kinetic_2020]. The bottom-right panel shows the geometric mean of measurements and predictions across conditions.
@@
Supplementary Figure 10: Turnover parameters used in EC-iCH360 before and after the adjustment procedure (see section 2.5 and Methods in the main text, as well as Supplementary Information A.5 for details). The original parameter set was directly parsed from [heckmann_kinetic_2020].
@@
Supplementary Figure 11: Impact of the magnitude of the ridge regularisation coefficient ($\rho$ in Eq. 35) on the outcome of the turnover number adjustment procedure. The blue curve represents the RMSE between measured and predicted log-enzyme abundances and decreases monotonically as less less regularisation is applied to the problem. The red curve represents the mean absolute deviation between original and adjusted log-turnover numbers, which follows the opposite trend and converges to 00 as more and more regularisation is applied.
@@
Supplementary Figure 12: Comparison of production envelopes for growth on glucose between iCH360 and the reduced variant iCH360red, used for elementary flux mode enumeration and analysis. The two models have virtually identical solution spaces for the products and growth conditions shown.
@@
Supplementary Figure 13: Flux distribution in the network corresponding to the maximum yield elementary flux mode (see main text, section 2.6). The mode is purely respiratory, with no excretion of typical fermentation by-products such as acetate, ethanol, or lactate. The graphics was produced in Escher [king_escher_2015].
@@
Supplementary Figure 14: Flux distribution in the network corresponding to the maximum growth elementary flux mode (see main text, section 2.6). The mode shows a mixed respiratory/fermentative phenotype, with significant acetate excretion, in line with typical observations of overflow metabolism. The graphics was produced in Escher [king_escher_2015].
@@
Supplementary Figure 15: EFM growth-yield tradeoff front under simulated low oxygen conditions, obtained by setting 1000-fold higher cost for the cytochrome reactions. The higher enzyme cost is used to mimic the growth condition in which the oxygen-consuming reaction must operate at very low enzyme saturation, thus requiring higher enzyme-mass investment per unit flux. Black squares denote Pareto optimal EFMs. The pareto front is much broader than the one obtained with the original enzyme cost for the oxygen-consuming reactions (see Figure 5 in main text).
@@
Supplementary Figure 16: satFBA results (corresponding to Figure 6 in the main text) obtained by enforcing a lower bound on the ATP maintenance flux equal to $6.86$ mmol/gDW/h (taken directly from the parent model iML1515). In this case, the optimal solution to the EFM is no longer an EFM, as evident from the yield profile (C), which is not piecewise constant with respect to the external glucose concentration.
@@
Supplementary Figure 17: Empirical cumulative distribution functions (CDF) of measured enzyme abundances for reactions with predicted flux force efficacy above or below $50\%$ (blue and red curves, respectively). The two are significantly different ($p<0.001$, two-sided Wilcox rank-sum test), with the low-efficacy group being associated to an approximately $3$-fold higher median median enzyme abundance than the high efficacy group.
@@
Supplementary Figure 18: Compressed metabolic map of iCH360, wherein linear pathways longer than two reactions were lumped together. The map was produced in Escher [king_escher_2015] and can thus be use to visualise fluxes, metabolite concentration, and gene expression data.

### v3

@@ A.5.5 Conversion to apparent turnover numbers
The above procedure produces (after conversion back to a linear scale) a set of adjusted turnover parameters, $\mathbf{k}^{*}$ and scalings, $\bm{\theta}^{*}$. In order to retrieve the typical enzyme efficiencies and condition-specific scaling factors introduced in 24, we simply parametrise the solution by factorising the scaling terms $\theta_{j}^{*}$ as
$\theta_{j}^{*}\equiv\bar{\sigma}^{*}~{}\tau_{j}^{*}$ (35)
Here, $\bar{\sigma}^{*}$ is the geometric mean of the scalings across conditions, which we interpret as typical enzyme saturation level across conditions, while $\tau_{j}^{*}$ is a residual scaling factor fluctuating around $1$ which is required to account for differences in total measured enzymes between conditions. The typical enzyme efficiencies ($\kappa_{i}$) introduced in 24 are then simply recovered by incorporating the $\bar{\sigma}^{*}$ constant into the fitted turnover parameters:
@@ A.5.6 Potential extensions
We conclude this section by noting that the procedure described above assumes that the original parameter set – the set of $k_{\rm cat}$ values used as proxies for apparent $k_{\rm cat}$ values in section A.5.3 – is sufficiently good to produce a realistic flux distribution to use as a reference for the adjustment step. If this is not the case, then a multi-start approach can be implemented, where multiple turnover parameter sets are first generated by perturbing the original parameter set, and each of them is used to generate a separate set of flux distributions. Each reference set is then provided as an input to problem (34), and the solution achieving the lowest objective is chosen in the end. The perturbed parameter set may be generated either randomly (for example, by introducing log-normal noise in the original turnover parameter vector) or systematically. The latter could be achieved, for example, by identifying reactions with zero-predicted flux but high measured abundance of the associated enzyme. By systematically increasing the corresponding turnover number, one can ”encourage” these reactions to be included in the reference set, potentially leading to the exploration of more relevant reference distributions than achieved by random perturbations.
@@ Appendix B Supplementary Figures
Supplementary Figure S1: Biosynthesis of amino acids (blue labels) from core metabolism precursors (red labels) in iCH360. The Escher maps (including Suppl. Fig. S2 to S4) are available along with the code. R5P: ribose-5-phosphate; E4P: erythrose 4-phosphate; PEP:phosphoenolpyruvate; 3PG: 3-phosphoglycerate; AKG: apha-ketoglutarate; CBP: carboamyl phosphate.
@@
Supplementary Figure S2: Biosynthesis of pyrimidine and purine (deoxy)ribonucleotides (blue labels) from core and aminoacid metabolism precursors (red labels) in iCH360. PRPP: 5-phosphoribosyl-1-pyrophosphate.
@@
Supplementary Figure S3: Biosynthesis of saturated and unsaturated fatty acids from acetyl-CoA. The map for saturated fatty acids was taken from Escher [king_escher_2015].
Supplementary Figure S4: Metabolism of one-carbon compounds in iCH360. THF: tetrahydrofolate.
@@
Supplementary Figure S5: Comparison of model sizes between ECC, ECC2, iCH360 and iML1515. To allow for a fair comparison, pseudo-reactions (e.g. exchange reactions) were excluded from the count. Note that gene annotations were not available in the SBML model of ECC2 accompanying its publication [hadicke_ecolicore2_2017].
@@
Supplementary Figure S6: Maximal biomass fluxes achieved by iCH360 and its parent model iML1515, for aerobic and anaerobic growth across multiple carbon sources. In all cases, the substrate uptake flux was bounded to $10$ mmol/gDW/h.
Supplementary Figure S7: Analysis of the differences in acetate production between iCH360 and its genome-scale parent, iML1515. A: available metabolic routes for the production of acetate in both models. Left: iCH360 can only produce acetyl-CoA, precursor for acetate, via the oxidation of pyruvate by either pyruvate dehydrogenase (PDH) or pyruvate-formate-lyase (PFL). Note that the latter reaction is known not to be active under aerobic conditions, but we did not block it for the purpose of this analysis. iML1515 can additionally produce acetate via additional pathways not present in iCH360 (in red). Note that only the main substrate and products for each reaction are shown for clarity. 2AOBUT: L-2-Amino-3-oxobutanoate; 2DR5P: 2-Deoxy-D-ribose 5-phosphate. B: Blocking these degradation routes by simultaneous knockout of four reactions (DRPA, PAI2T, THRA, THRD) results in the two model sharing a virtually identical production envelope under aerobic conditions. The production envelope shown here was computed for aerobic growth using glucose as a carbon source. C: Under anaerobic conditions, the differences between the two model are further exacerbated by the ability of the genome-scale network to achieve higher pyruvate production yield. The genome-scale model can produce higher amounts of pyruvate by uptaking external CO2 and using it as an electron sink using the POR5 reaction. FLX ox/red: oxidised/reduced flavodoxin. D: Blocking CO2 uptake reduces the differences in pyruvate production between the two models, but does not remove them completely, implying the existence of additional mechanisms used by iML1515 to achieve higher pyruvate yields. The production envelopes shown here was computed for anaerobic growth using glucose as a carbon source.
@@
Supplementary Figure S8: Comparison of production envelopes between iCH360 and its parent model iML1515. Top: production of ethanol, acetate and succinate during aerobic growth on lactate. Bottom: production of ethanol, acetate, lactate and succinate during aerobic growth on glycerol. Note that the dashed line representing the production envelope of iML1515 is sometimes hidden behind the blue line corresponding to iCH360.
Supplementary Figure S9: Comparison of production envelopes between iCH360 and other medium-scale models, namely E. coli Core (ECC) and E. coli Core 2 (ECC2) for growth on glucose as a sole carbon source.Top: production of ethanol, acetate and succinate under aerobic conditions. Bottom: production of ethanol, acetate, lactate and succinate under anaerobic conditions. Additional comparisons between the three models are available in the repository supporting this manuscript.
@@
Supplementary Figure S10: Fitness losses associated with disruptions of catalytic edges. There is no significant difference between the fitness effects of disruptions classified as complete disruptions (disruption of all catalytic edges for a reaction) and full primary disruption (disruption of all primary catalytic associations, but with remaining secondary ones), according to a Wilcoxon rank-sum test, $p=0.063$.
@@
Supplementary Figure S11: Predicted proteome allocation across growth conditions, obtained with the turnover parameter set from [heckmann_kinetic_2020]. The bottom-right panel shows the geometric means of measurements and predictions across conditions.
@@
Supplementary Figure S12: Turnover parameters used in EC-iCH360 before and after the adjustment procedure (see section 2.4 and Methods in the main text, as well as Supplementary Information A.5 for details). The original parameter set was directly parsed from [heckmann_kinetic_2020].
@@
Supplementary Figure S13: Impact of the magnitude of the ridge regularisation coefficient ($\rho$ in Eq. (34)) on the outcome of the adjustment procedure for turnover numbers. The blue curve represents the RMSE between measured and predicted log-enzyme abundances and decreases monotonically as less regularisation is applied to the problem. The red curve represents the mean absolute deviation between original and adjusted log-turnover numbers, which follows the opposite trend and converges to 00 as more regularisation is applied.
Supplementary Figure S14: Leave-one-out cross-validation for the turnover parameter adjustment procedure. For each condition in the dataset, the graph shows the RMSE (computed for log10-transformed values) between measurements and predictions of enzyme abundances in that condition. Blue bars show the RMSE computed using the initial, unadjusted parameter set from [heckmann_kinetic_2020]. Red bars show the RMSE computed after parameter fitting, but excluding the condition for which the RMSE is evaluated from the training dataset. Finally, yellow bars show the RMSE computed using the final adjusted parameter set, obtained including all conditions in the training dataset.
@@
Supplementary Figure S15: Comparison of production envelopes for growth on glucose between iCH360 and the reduced variant iCH360red used for elementary flux mode enumeration and analysis. For the products and growth conditions shown, the two models have virtually identical solution spaces.
@@
Supplementary Figure S16: Metabolic flux distribution corresponding to the maximum yield elementary flux mode (see main text, section 2.5). The mode is purely respiratory, with no excretion of typical fermentation by-products such as acetate, ethanol, or lactate. The graphics was produced in Escher [king_escher_2015].
@@
Supplementary Figure S17: Metabolic flux distribution corresponding to the maximum growth elementary flux mode (see main text, section 2.5). The mode shows a mixed respiratory/fermentative phenotype, with significant acetate excretion. The graphics was produced in Escher [king_escher_2015].
@@
Supplementary Figure S18: EFM growth-yield trade-off front under simulated low oxygen conditions. The figure resembles Figure 5 in the main text, but was obtained by setting a 1000-fold higher cost for the cytochrome reactions. The higher enzyme cost is used to mimic the growth condition in which the oxygen-consuming reaction must operate at very low enzyme saturation, thus requiring higher enzyme-mass investment per unit flux. Black squares denote Pareto optimal EFMs. The Pareto front is much broader than the one obtained with the original enzyme cost for the oxygen-consuming reactions.
@@
Supplementary Figure S19: satFBA results (corresponding to Figure 6 in the main text) obtained by enforcing a lower bound on the ATP maintenance flux equal to $6.86$ mmol/gDW/h (taken directly from the parent model iML1515). In this case, the optimal solution is no longer an EFM. This is evident from the yield profile (C), which is not piecewise constant with respect to the external glucose concentration.
@@
Supplementary Figure S20: Empirical cumulative distribution functions (CDF) of measured enzyme abundances for reactions with predicted flux force efficacies above or below $50\%$ (blue and red curves, respectively). The two functions are significantly different ($p<0.001$, two-sided Wilcox rank-sum test). The low-efficacy group shows an approximately $3$-fold higher median enzyme abundance than the high efficacy group.
@@
Supplementary Figure S21: Compressed metabolic map of iCH360, wherein linear pathways longer than two reactions were lumped into single effective reactions. The map was produced in Escher [king_escher_2015] and can be used to visualise fluxes, metabolite concentration, and gene expression data.
Supplementary Figure S22: Computation of boolean gene-protein-reaction (GPR) rules with the knowledge graph. For each node in the graph (the top node in each diagram), a boolean expression can be constructed to describe the state of a node (active or inactive) in terms of the state of its child nodes. The exact form of this boolean expression depends both on the node type (e.g. protein, reaction, or logical) and on the type of edges that connect it with its neighbours (e.g. catalysis, subunit composition, etc.). Using these rules, the state of a reaction can be expressed, ultimately, solely in terms of genes in the model, as per convention in standard metabolic models. These GPR rules define a map between a genotype (set of active genes) to a phenotype (set of active reactions, and can be used to simulate in silico the effect of given gene knockouts.
@@ Appendix C Supplementary Tables
Supplementary Table 1: Some anecdotal examples of unrealistic predictions obtained when running FBA analyses on large-scale models (here, we considered the genome-scale parent of iCH360, iML1515). We stress that these behaviours are not the result of errors in the large model. Rather, they result from applying simple methods with few constraints, such as FBA, to a large model with many degrees of freedom. Note that the examples shown here were not obtained through a thorough or systematic investigation of the model prediction abilities.
[…]
@@
(In this example, the logical edges are used to indicate that the THIOREDOXINS node (of type logical-OR) is active when any of the two child nodes (representing two thioredoxins found in E. coli) are active.